Faced with headquarters’ desire to add a new product line, Stefan Grenier, manager of Bilti Products’ East Division, felt that he had to see the numbers before he made a move. His division’s ROI has led the company for three years, and he doesn’t want any letdown.

     Bilti Products is a decentralized wholesaler with four autonomous divisions. The divisions are evaluated on the basis of ROI, with year-end bonuses given to divisional managers who have the highest ROI. Operating results for the company’s East Division for last year are given below:

  Sales $ 34,300,000
  Variable expenses 15,110,000
  
  Contribution margin 19,190,000
  Fixed expenses 16,446,000
  Operating income $ 2,744,000
  Divisional operating assets $ 8,575,000


The company had an overall ROI of 16% last year (considering all divisions). The new product line that headquarters wants Grenier’s East Division to add would require an investment of $4,900,000. The cost and revenue characteristics of the new product line per year would be as follows:

  Sales $ 14,700,000
  Variable expenses 70% of sales
  Fixed expenses $ 3,528,000


Required:
1. Compute the East Division’s ROI for last year; also compute the ROI as it would appear if the new product line were added. (Do not round intermediate calculations. )

2. If you were in Grenier’s position, would you accept or reject the new product line?

  • Accept

  • Reject

3. Why do you suppose headquarters is anxious for the East Division to add the new product line?

  • Adding the new line would increase the company's overall ROI.

  • Adding the new line would decrease the company's overall ROI.

4. Suppose that the company’s minimum required rate of return on operating assets is 15% and that performance is evaluated using residual income.

a. Compute East Division’s residual income for last year; also compute the residual income as it would appear if the new product line were added.

b. Under these circumstances, if you were in Grenier’s position, would you accept or reject the new product line?

  • Accept

  • Reject
Answer #1

1. ROI = Operating Income/Operating Assets

= 2,744,000/8,575,000

= 32%

If new line added, ROI = (2,744,000+14,700,000*30%-3,528,000)/(8,575,000+4,900,000)

= 26.91%

Reject, since it will reduce ROI

3. Adding the new line would increase the company's overall ROI.

a.Residual Income = Operating Income – Operating Assets*Required rate of return

= 2,744,000-8,575,000*15%

= $1,457,750

If Added, RI = 1,457,750 + 882,000 – 4,900,000*15%

= $1,604,750

Accept, since it will increase Residual Income